# Convolution estimates for the Boltzmann gain operator with hard-spheres

### v3

@@ 1.2. Heuristics
However, despite the point-wise growth, we show that these singular energy-absorbing collisions do not drastically affect the averaging behavior of the gain operator. In other words, moment-preserving estimates still hold for the gain operator. The key idea is that, although energy-absorbing collisions occur with non-trivial probability, they are statistically rare. This is quantified by a point-wise cancellation and singular angular integration, see Lemma 2.5 (Cancellation Lemma) below. This estimate plays a crucial role in our analysis.
We now explain in more detail how these energy-absorbing collisions arise. They constitute the most serious obstruction to proving moment-preserving estimates.
Large velocity splitting and convolution estimates
In order to estimate $Q^{+}$, we first decompose it into two pieces $Q_{0},Q_{1}$ corresponding to the outgoing particle with the largest velocity. Namely, by conservation of energy we have either $|v^{*}|^{2}\geq E/2$ or $|v_{1}^{*}|^{2}\geq E/2$, thus we can bound
(1.15) $|Q^{+}(f,g)|\lesssim|Q_{0}(f,g)|+|Q_{1}(f,g)|,$
where
(1.16) $Q_{0}(f,g)$ $=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|u|f(v^{*})g(v_{1}^{*})\chi(\widehat{u}\cdot\sigma)\mathds{1}_{|v^{*}|^{2}\geq E/2}\,d\sigma\,dv_{1},$
(1.17) $Q_{1}(f,g)$ $=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|u|f(v^{*})g(v_{1}^{*})\chi(\widehat{u}\cdot\sigma)\mathds{1}_{|v_{1}^{*}|^{2}\geq E/2}\,d\sigma\,dv_{1}.$
Therefore it suffices to prove convolution estimates for $Q_{0}$ and $Q_{1}$. This is the content of Propositions 4.14.4, which is the heart of our contribution. Upon establishing those estimates, the claim of Theorem 1.1 follows easily, see Subsection 4.5.
The main difficulty in this task is that the operators $Q_{0}$ and $Q_{1}$ are singular (in the sense that $\mathcal{R}\approx|v|$) in the regions $v_{1}^{*}\approx 0$ and $v^{*}\approx 0$ respectively, which corresponds exactly to the energy-absorbing collisions mentioned above.
As we will see, $Q_{0}$ exhibits more regular behavior due to the cut-off $\chi$ excluding some very pathological interactions. $Q_{1}$ on the other hand is more intricate since these interactions are present. Roughly speaking, $Q_{0}$ has a simple singularity, while $Q_{1}$ has a double singularity.
More specifically, the cut-off $\chi$ prevents interactions of the type $v_{1}^{*}\approx 0$ and $v_{1}\approx 0$ at the same time. To see this formally, assume $|v_{1}|\ll|v|$. Then $u=v-v_{1}\approx v$, so $|v_{1}|\ll|u|$. Using (1.3), we can write
$v_{1}^{*}=v_{1}+R_{\sigma}^{+}(u),\quad R_{\sigma}^{+}(u):=\frac{u}{2}+\frac{|u|}{2}\sigma.$
Now, the cut-off $\chi$ restricts us to the region where $\widehat{u}\cdot\sigma>0$, hence
$|R_{\sigma}^{+}(u)|^{2}=\frac{|u|^{2}}{2}\left(1+\widehat{u}\cdot\sigma\right)>\frac{|u|^{2}}{2}.$
By the triangle inequality, we obtain
$|v_{1}^{*}|\geq|R_{\sigma}^{+}(u)|-|v_{1}|>\frac{|u|}{\sqrt{2}}-|v_{1}|>\frac{|u|}{2}\gtrsim|v|,$
since $|v_{1}|\ll|u|\approx|v|$. In other words, we cannot have $v_{1}\approx 0$ and $v_{1}^{*}\approx 0$ simultaneously; therefore we can always assume that $|v_{1}|\gtrsim|v|$ when approximating the singularity of $Q_{0}$.
This reasoning does not apply to the singularity of $Q_{1}$ since
$v^{*}=v_{1}+R_{\sigma}^{-}(u),\quad R_{\sigma}^{-}(u):=\frac{u}{2}-\frac{|u|}{2}\sigma,$
and
$|R_{\sigma}^{-}(u)|^{2}=\frac{|u|^{2}}{2}\left(1-\widehat{u}\cdot\sigma\right).$
Since $\widehat{u}\cdot\sigma$ is allowed to be arbitrarily close to $1$, we can have $R_{\sigma}^{-}(u)\approx 0$, and thus $v^{*}\approx v_{1}\approx 0$. This visually corresponds to a full transfer of energy: a particle carries all the energy before the collision, it collides with a still particle, and the energy is fully transferred to the previously still particle. This situation is the most pathologic one, and results in an additional singularity which makes the estimate of $Q_{1}$ more challenging.
@@ 1.3. Main results
We now present the main results of this paper more formally. Given $k\geq 0$ and $1\leq p,q\leq\infty$, we define the normed space
@@
Note that the convergence of a sequence $\{f_{n}\}_{n=1}^{\infty}\subset\bm{\mathcal{X}_{p,q}^{k}}$ to some $f\in\bm{\mathcal{X}_{p,q}^{k}}$ can be characterized as follows:
@@
Let $0<\gamma\leq 1$, $b$ satisfying (H.1)–(H.3), $1<p<2<q<\infty$, $2<r\leq\infty$ such that $\frac{1}{p}+\frac{1}{q}=1+\frac{1}{r}$ and $k\geq l>2$. Then, $Q^{+}$ extends uniquely to a bounded, bilinear operator $Q^{+}:\bm{\mathcal{X}_{p,q}^{k}}\times\bm{\mathcal{X}_{p,q}^{k}}\to{\langle}v{\rangle}^{-k}L^{r}_{v}$ which satisfies the estimate
(1.21) $\|{\langle}v{\rangle}^{k}Q^{+}(f,g)\|_{L_{v}^{r}}\lesssim\|{\langle}v{\rangle}^{k}f\|_{L_{v}^{p}}\|{\langle}v{\rangle}^{l}g\|_{L_{v}^{q}}+\|{\langle}v{\rangle}^{l}f\|_{L_{v}^{p}}\|{\langle}v{\rangle}^{k}g\|_{L_{v}^{q}}+\|{\langle}v{\rangle}^{l}f\|_{L_{v}^{q}}\|{\langle}v{\rangle}^{k}g\|_{L_{v}^{p}},$
for all $f,g\in\bm{\mathcal{X}_{p,q}^{k}}$.
@@
Unlike for polynomial weights, we do not expect moment-preserving bounds on $Q^{+}$ to hold. Indeed, consider Gaussians of the form $e^{|v|^{2}}$. To prove the corresponding weighted estimates, one would need to control the ratio
@@
  • In Section 2, we provide the necessary kinetic tools that will be used in the rest of the paper. Namely, we introduce the Bobylev variables and prove Lemma 2.2 and Lemma 2.5 which are instrumental in our analysis.
  • In Section 3, we prove convolution estimates for Maxwell molecules. Theorem 3.1 provides the convolution estimate for $r=\infty$, while Theorem 3.2 provides the convolution estimate for $r<\infty$.
  • In Section 4, we prove the convolution estimates for hard potentials, which is our most noteworthy contribution. Propositions 4.1, 4.3 provide the estimates for $Q_{0}$, while Propositions 4.2, 4.4 provide the estimates for $Q_{1}$. We conclude with the proof of Theorem 1.1.
@@
We will also be using the following notation:
  • We write $\chi:=\mathds{1}_{(0,+\infty)}$ for the standard cut-off function.
  • We write $C^{\infty}_{c}({\mathbb{R}}^{3})$ for the set of smooth, compactly supported functions in ${\mathbb{R}}^{3}$.
@@ Acknowledgments
I.A. was supported by the NSF grant No. DMS-2418020 and the PSC-CUNY Research Award 68653-00 56. The authors are also grateful to the Institute for Advanced Study for providing a peaceful environment to conduct this research during July 2026.
@@ 2. Technical toolbox
In this section, we present some technical results fundamental to our analysis. First, we introduce the so-called Bobylev variables [9, 10], and then employ them to prove key angular averaging estimates. Namely, Lemma 2.2 will be used to treat Maxwell molecules, while Lemma 2.5 unveils the cancellation property of the gain collisional operator which allows us to control energy-absorbing collisions, and therefore treat hard potentials. Finally, we record the well-known pre-post collisional change of variables in Lemma 2.6.
@@ 2.1. Bobylev variables
We now state some finer substitution properties of $R_{\sigma}^{+},R_{\sigma}^{-}$ that will be important in uncovering the collisional averaging behavior of the gain operator. The proof of the following result can be found in e.g. the work of Arsenio [7], as well as our previous work [3]. For convenience of the reader though, we present a detailed proof.
@@
(2.10) $|R_{\sigma}^{\epsilon}(y)\cdot\sigma|$ $=\epsilon\,(R_{\sigma}^{\epsilon}(y\cdot\sigma),$
@@
thus $R_{\sigma}^{\epsilon}:A_{\sigma}^{\epsilon}\to B_{\sigma}^{\epsilon}$, as well as (2.10) holds.
@@
2.2. Maxwell collisional averaging
Here, we use Lemma 2.1 to obtain the main averaging estimates for Maxwell molecules:
@@ 2.2. Maxwell collisional averaging
Let $\beta_{1}<1$, $\beta_{2}>1/2$ and $1\leq p<\infty$. For $f\in C_{c}^{\infty}({\mathbb{R}}^{3})$, we denote
(2.16) $\mathcal{I}^{1}_{\beta_{1},p}[f](v)$ $:=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\frac{|f(v^{*})|^{p}\chi(\widehat{u}\cdot\sigma)}{|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2\beta_{1}}}\,d\sigma\,dv_{1},$
@@
Consider $f\in C_{c}^{\infty}({\mathbb{R}}^{3})$. We first prove (2.18). Fix $v\in{\mathbb{R}}^{3}$. Using (2.6), (2.12), (2.13) and the substitution $y:=u=v-v_{1}$ we write
@@ 2.3. Cancellation Lemma
Now, we use Lemma 2.1 to prove Lemma 2.5 which encodes the cancellation mechanism of the gain collisional operator. First, we present an important point-wise cancellation estimate in Lemma 2.3 and a singular integral estimate in Lemma 2.4:
@@
For any $v,v_{1}\in{\mathbb{R}}^{3}$, $\sigma\in{\mathbb{S}}^{2}$, the following bounds holds
(2.20) ${\langle}v^{*}{\rangle}$ $\geq\frac{{\langle}v_{1}^{*}{\rangle}}{\sqrt{2}}\left(1-\lambda_{v_{1}^{*}}|\widehat{v_{1}^{*}}\cdot\sigma|^{2}\right)^{1/2},\quad\lambda_{v_{1}^{*}}=\frac{|v_{1}^{*}|^{2}}{{\langle}v_{1}^{*}{\rangle}^{2}},$
(2.21) ${\langle}v_{1}^{*}{\rangle}$ $\geq\frac{{\langle}v^{*}{\rangle}}{\sqrt{2}}\left(1-\lambda_{v^{*}}|\widehat{v^{*}}\cdot\sigma|^{2}\right)^{1/2},\quad\lambda_{v^{*}}=\frac{|v^{*}|^{2}}{{\langle}v^{*}{\rangle}^{2}}.$
@@
Fix $v,v_{1}\in{\mathbb{R}}^{3}$ and $\sigma\in{\mathbb{S}}^{2}$. Using the elementary inequality
$|u||V|\leq\frac{|u|^{2}}{4}+|V|^{2}=\frac{E}{2},$
we obtain
$|v^{*}|^{2}$ $=\frac{E}{2}-|u||V|(\widehat{V}\cdot\sigma)\geq\frac{E}{2}(1-|\widehat{V}\cdot\sigma|)=\frac{E}{2}\frac{1-|\widehat{V}\cdot\sigma|^{2}}{1+|\widehat{V}\cdot\sigma|}\geq\frac{E}{4}(1-|\widehat{V}\cdot\sigma|^{2})$
$\geq\frac{|V|^{2}}{2}(1-|\widehat{V}\cdot\sigma|^{2}).$
Now, using the expression $V=v_{1}^{*}-\frac{|u|}{2}\sigma$, we compute
$|V|^{2}(1-|\widehat{V}\cdot\sigma|^{2})$ $=\left|v_{1}^{*}-\frac{|u|}{2}\sigma\right|^{2}-\left|v_{1}^{*}\cdot\sigma-\frac{|u|}{2}\right|^{2}=|v_{1}^{*}|^{2}(1-|\widehat{v_{1}^{*}}\cdot\sigma|^{2}),$
thus $|v^{*}|^{2}\geq\frac{|v_{1}^{*}|^{2}}{2}(1-|\widehat{v_{1}^{*}}\cdot\sigma|^{2})$. Hence, we obtain
${\langle}v_{1}^{*}{\rangle}^{2}$ $=1+|v_{1}^{*}|^{2}\geq\frac{1}{2}\left(1+|v_{1}^{*}|^{2}(1-|\widehat{v_{1}^{*}}\cdot\sigma|^{2})\right)=\frac{{\langle}v_{1}^{*}{\rangle}^{2}}{2}(1-\lambda_{v_{1}^{*}}|\widehat{v_{1}^{*}}\cdot\sigma|^{2}),$
where $\lambda=\frac{|v_{1}^{*}|^{2}}{{\langle}v_{1}^{*}{\rangle}^{2}}$, and (2.20) is proved.
The bound (2.21) is proved identically using the expression $V=v^{*}+\frac{|u|}{2}\sigma$ instead. ∎
@@
(2.22) $\int_{{\mathbb{S}}^{2}}\frac{1}{(1-\lambda|\widehat{\eta}\cdot\sigma|^{2})^{l/2}}\,d\sigma\lesssim\frac{1}{(l-2)\lambda}(1-\lambda)^{1-l/2}.$
@@
$\int_{{\mathbb{S}}^{2}}\frac{1}{(1-\lambda|\widehat{\eta}\cdot\sigma|^{2})^{l/2}}\,d\sigma$ $\approx\int_{0}^{1}\frac{1}{(1-\lambda x^{2})^{l/2}}\,dx\leq\int_{0}^{1}\frac{1}{(1-\lambda x)^{l/2}}\,dx$
$=\frac{1}{\lambda}\int_{0}^{\lambda}\frac{1}{(1-z)^{l/2}}\,dz\leq\frac{1}{(l-2)\lambda}(1-\lambda)^{1-l/2},$
@@
We are now in the position to prove Lemma 2.5:
@@
Let $l>2$ and $1\leq p<\infty$. For $f\in C_{c}^{\infty}({\mathbb{R}}^{3})$, we denote
(2.23) $\mathcal{J}_{l,p}^{1}[f](v)$ $:=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\frac{|f(v^{*})|^{p}\chi(\widehat{u}\cdot\sigma)}{{\langle}v_{1}^{*}{\rangle}^{l}}\mathds{1}_{|v|>4}\mathds{1}_{|v^{*}|^{2}\geq E/2}\,d\sigma\,dv_{1},$
(2.24) $\mathcal{J}_{l,p}^{2}[f](v)$ $:=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\frac{|f(v_{1}^{*})||\widehat{R}_{\sigma}^{-}(u)|^{2}}{{\langle}v^{*}{\rangle}^{l}}\mathds{1}_{|v|>4}\mathds{1}_{|v_{1}^{*}|^{2}\geq E/2}\,d\sigma\,dv_{1}.$
Then there holds the estimates
(2.25) $\|{\langle}v{\rangle}^{2}\mathcal{J}_{l,p}^{1}[f]\|_{L_{v}^{\infty}}$ $\lesssim\|f\|_{L_{v}^{p}}^{p},$
(2.26) $\|{\langle}v{\rangle}^{2}\mathcal{J}_{l,p}^{2}[f]\|_{L_{v}^{\infty}}$ $\lesssim\|f\|_{L_{v}^{p}}^{p}.$
@@
Throughout the proof, given $a\in{\mathbb{R}}^{3}$ with $a\neq 0$, we will denote $\lambda_{a}:=\frac{|a|^{2}}{{\langle}a{\rangle}^{2}}\in(0,1)$. Consider $f\in C_{c}^{\infty}({\mathbb{R}}^{3})$.
Estimate for $\mathcal{J}_{l,p}^{1}$
We first prove (2.25). Fix $v\in{\mathbb{R}}^{3}$. We use (2.21), followed by (2.6) and (2.12) to bound
$\mathcal{J}_{l,p}^{1}[f](v)\lesssim\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\frac{|f(v^{*})|^{p}\chi(\widehat{u}\cdot\sigma)}{{\langle}v^{*}{\rangle}^{l}\left(1-\lambda_{v^{*}}|\widehat{v^{*}}\cdot\sigma|^{2}\right)^{l/2}}\mathds{1}_{|v^{*}|>|v|/2>2}\,d\sigma\,dv_{1}$
$=\int_{{\mathbb{S}}^{2}}\int_{{\mathbb{R}}^{3}}\frac{|f(v-R_{\sigma}^{+}(u))|^{p}\chi(2|\widehat{R}^{+}_{\sigma}(u)\cdot\sigma|^{2}-1)}{\left\langle v-R_{\sigma}^{+}(u)\right\rangle^{l}\left(1-\lambda_{v-R_{\sigma}^{+}(u)}|\widehat{(v-R_{\sigma}^{+}(u))}\cdot\sigma|^{2}\right)^{l/2}}\mathds{1}_{|v-R_{\sigma}^{+}(u)|>|v|/2>2}\,dv_{1}\,d\sigma$
$=\int_{{\mathbb{S}}^{2}}\int_{{\mathbb{R}}^{3}}\frac{|f(v-R_{\sigma}^{+}(y))|^{p}\chi(2|\widehat{R}^{+}_{\sigma}(y)\cdot\sigma|^{2}-1)}{{\langle}v-R_{\sigma}^{+}(y){\rangle}^{l}\left(1-\lambda_{v-R_{\sigma}^{+}(y)}|\widehat{(v-R_{\sigma}^{+}(y))}\cdot\sigma|^{2}\right)^{l/2}}\mathds{1}_{|v-R_{\sigma}^{+}(y)|>|v|/2>2}\,dy\,d\sigma.$
Using Proposition 2.1 to substitute $\nu:=R_{\sigma}^{+}(y)$, we obtain
$\mathcal{J}_{l,p}^{1}[f](v)\lesssim\int_{{\mathbb{S}}^{2}}\int_{{\mathbb{R}}^{3}}{\langle}v-\nu{\rangle}^{-l}|f(v-\nu)|^{p}\mathds{1}_{|v-\nu|>\frac{|v|}{2}>2}\frac{\chi(2|\widehat{\nu}\cdot\sigma|^{2}-1)}{|\widehat{\nu}\cdot\sigma|^{2}\left(1-\lambda_{v-\nu}|\widehat{(v-\nu)}\cdot\sigma|^{2}\right)^{l/2}}\,d\nu\,d\sigma$
$\approx\int_{{\mathbb{R}}^{3}}{\langle}v-\nu{\rangle}^{-l}|f(v-\nu)|^{p}\mathds{1}_{|v-\nu|>\frac{|v|}{2}>2}\int_{{\mathbb{S}}^{2}}\frac{1}{\left(1-\lambda_{v-\nu}|\widehat{(v-\nu)}\cdot\sigma|^{2}\right)^{l/2}}\,d\sigma\,d\nu,$
where we used the basic inequality $\frac{\chi(2z^{2}-1)}{z^{2}}\lesssim 1$, $z\in[0,1]$.

### v4

@@ 1.2. Heuristics
However, despite the point-wise growth, we show that these singular energy-absorbing collisions do not drastically affect the averaging behavior of the gain operator. In other words, moment-preserving estimates still hold for the gain operator. The key idea is that, although energy-absorbing collisions occur with non-trivial probability, they are statistically rare. This is quantified by a cancellation and singular angular integration, see Lemma 2.5 (Cancellation Lemma) below. This estimate plays a crucial role in our analysis.
@@ 1.3. Main results
We now present the main results of this paper more formally. Given $k\geq 0$ and $1\leq p,q\leq\infty$, we define the space
@@
Note that the convergence of a sequence $\{f_{n}\}_{n=1}^{\infty}\subset\bm{\mathcal{X}_{p,q}^{k}}$ to some $f\in\bm{\mathcal{X}_{p,q}^{k}}$ can be equivalently characterized as follows:
@@
Let $0<\gamma\leq 1$ and $b$ satisfying (H.1)–(H.3). Consider $1<p<2<q<\infty$, $2<r\leq\infty$ such that $\frac{1}{p}+\frac{1}{q}=1+\frac{1}{r}$, and $k\geq l>1$ with $k\geq 2+\gamma-\frac{2}{q}$, $l>2-\frac{2}{q}$. Then, $Q^{+}$ extends uniquely to a bounded, bilinear operator $Q^{+}:\bm{\mathcal{X}_{p,q}^{k}}\times\bm{\mathcal{X}_{p,q}^{k}}\to{\langle}v{\rangle}^{-k-1+\gamma}L^{r}_{v}$ which satisfies the estimate
(1.19) $\|{\langle}v{\rangle}^{k+1-\gamma}Q^{+}(f,g)\|_{L_{v}^{r}}\lesssim\|{\langle}v{\rangle}^{l}f\|_{L_{v}^{p}}\|{\langle}v{\rangle}^{k}g\|_{L_{v}^{q}}+\|{\langle}v{\rangle}^{l}f\|_{L_{v}^{q}}\|{\langle}v{\rangle}^{k}g\|_{L_{v}^{p}},\quad\forall\,f,g\in\bm{\mathcal{X}_{p,q}^{k}}.$
@@
The first term on the right-hand side of (1.19) roughly corresponds to the piece of the operator where $|v_{1}|\ll|v|$ (see Figure 3), while the second term corresponds to $|v_{1}|\gtrsim|v|$, see Figure 2. The relative sizes of these velocities will be carefully quantified in terms of $|\widehat{\nu}\cdot\sigma|$, where $\nu=v-v_{1}^{*}$. As we will see, the second term controls the operator as long as $|\widehat{\nu}\cdot\sigma|>{\langle}v{\rangle}^{-1/2}$, while the first term arises in the singular region where $|\widehat{\nu}\cdot\sigma|<{\langle}v{\rangle}^{-1/2}$.
Remark 1.3.
In the case of a hard potential less singular than hard-spheres i.e. $\gamma<1$, our estimates are in fact moment-decreasing.
Remark 1.4.
Unlike for polynomial weights, we do not expect moment-preserving bounds on $Q^{+}$ to hold. Indeed, consider Gaussian weights of the form $e^{|v|^{2}}$. To prove the corresponding weighted estimates, one would need to control the ratio
@@
  • In Section 2, we provide the necessary kinetic tools that will be used in the rest of the paper. Namely, we introduce the Bobylev variables, prove Lemma 2.2 and Lemma 2.5, which are instrumental in our analysis.
  • In Section 3, we prove the convolution estimates for Maxwell molecules. Theorem 3.1 provides the convolution estimate for $r=\infty$, while Theorem 3.2 provides the convolution estimate for $r<\infty$.
  • In Section 4, we prove Theorem 1.1 which provides moment-preserving convolution estimates for hard-spheres and is the heart of our contribution.
@@
We will also use the following notation:
  • We write $C_{c}({\mathbb{R}}^{3})$ for the space of continuous, compactly supported functions in ${\mathbb{R}}^{3}$, $C_{c}^{\infty}({\mathbb{R}}^{3})$ for the space of smooth, compactly supported functions and $C_{b}({\mathbb{R}}^{3})$ for the space of continuous and bounded functions.
@@ Acknowledgments
I.A. was supported by NSF grant No. DMS-2418020 and the PSC-CUNY Research Award 68653-00 56. The authors are also grateful to the Institute for Advanced Study for providing a peaceful environment to conduct this research during July 2025.
@@ 2. Technical toolbox
In this section, we present some technical results fundamental to our analysis. First, we introduce the so-called Bobylev variables [9, 10], and then employ them to prove key angular averaging estimates. Namely, Lemma 2.2 will be used to treat Maxwell molecules, while Lemma 2.5 captures the cancellation property of the gain collisional operator, which allows us to control energy-absorbing collisions and treat hard potentials. Finally, we record the well-known pre-post collisional change of variables in Lemma 2.6.
@@ 2.1. Bobylev variables
We now state some finer substitution properties of $R_{\sigma}^{+},R_{\sigma}^{-}$ that will be important in quantifying the collisional averaging behavior of the gain operator. The proof of the following result can be found in e.g. the work of Arsenio [7], as well as our previous work [3]. For convenience of the reader we present a detailed proof.
@@
(2.10) $|R_{\sigma}^{\epsilon}(y)\cdot\sigma|$ $=\epsilon\,(R_{\sigma}^{\epsilon}(y)\cdot\sigma),$
@@
thus $R_{\sigma}^{\epsilon}:A_{\sigma}^{\epsilon}\to B_{\sigma}^{\epsilon}$, and moreover (2.10) holds.
@@
2.2. Basic collisional averaging
Here, we use Lemma 2.1 to obtain the main averaging estimates when moments do not increase. Those estimates will be particularly useful for understanding Maxwell molecules, see Section 3.
@@ 2.2. Basic collisional averaging
Let $\beta_{1}<1$, $\beta_{2}>1/2$ and $1\leq p<\infty$. For $f\in C_{c}({\mathbb{R}}^{3})$, we denote
(2.16) $\mathcal{I}^{1}_{\beta_{1},p}[f](v)$ $:=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\frac{|f(v^{*})|^{p}\mathds{1}_{\widehat{u}\cdot\sigma>0}}{|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2\beta_{1}}}\,d\sigma\,dv_{1},$
@@
Consider $f\in C_{c}({\mathbb{R}}^{3})$. We first prove (2.18). Let us write $\chi(x):=\mathds{1}_{(0,\infty)}(x)$ and fix $v\in{\mathbb{R}}^{3}$. Using (2.6), (2.12), (2.13) and the substitution $y:=u=v-v_{1}$ we write
@@ 2.3. Cancellation Lemma
Now we use Lemma 2.1 to prove Lemma 2.5 which encodes the cancellation mechanism of the gain collisional operator for hard potentials. First, we present an important point-wise cancellation estimate in Lemma 2.3 and an elementary singular integral estimate in Lemma 2.4:
@@
For any $v,v_{1}\in{\mathbb{R}}^{3}$, $\sigma\in{\mathbb{S}}^{2}$, the following bounds hold
(2.20) ${\langle}v^{*}{\rangle}$ $\geq{\langle}v_{1}^{*}{\rangle}\left(1-\lambda_{v_{1}^{*}}|\widehat{v_{1}^{*}}\cdot\sigma|\right)^{1/2},\quad\lambda_{v_{1}^{*}}=\frac{|v_{1}^{*}|^{2}}{{\langle}v_{1}^{*}{\rangle}^{2}},$
(2.21) ${\langle}v_{1}^{*}{\rangle}$ $\geq{\langle}v^{*}{\rangle}\left(1-\lambda_{v^{*}}|\widehat{v^{*}}\cdot\sigma|\right)^{1/2},\quad\lambda_{v^{*}}=\frac{|v^{*}|^{2}}{{\langle}v^{*}{\rangle}^{2}}.$
@@
Fix $v,v_{1}\in{\mathbb{R}}^{3}$ and $\sigma\in{\mathbb{S}}^{2}$. By (1.3) we have $v^{*}=v_{1}^{*}-|u|\sigma$, so we compute
$|v^{*}|^{2}=|v_{1}^{*}|^{2}+|u|^{2}-2|u||v_{1}^{*}|(\widehat{v}_{1}^{*}\cdot\sigma)\geq(|v_{1}^{*}|^{2}+|u|^{2})\left(1-\frac{2|u||v_{1}^{*}|}{|v_{1}^{*}|^{2}+|u|^{2}}|\widehat{v_{1}^{*}}\cdot\sigma|\right)\geq|v_{1}^{*}|^{2}(1-|\widehat{v_{1}^{*}}\cdot\sigma|),$
where we used the basic inequality $2|u||v_{1}^{*}|\leq|u|^{2}+|v_{1}^{*}|^{2}$. Therefore
${\langle}v_{1}^{*}{\rangle}^{2}=1+|v_{1}^{*}|^{2}\geq(1+|v_{1}^{*}|^{2})\left(1-\frac{|v_{1}^{*}|^{2}}{1+|v_{1}^{*}|^{2}}|\widehat{v_{1}^{*}}\cdot\sigma|\right)={\langle}v_{1}^{*}{\rangle}^{2}\left(1-\lambda_{v_{1}^{*}}|\widehat{v_{1}^{*}}\cdot\sigma|\right),$
where $\lambda_{v_{1}^{*}}=\frac{|v_{1}^{*}|^{2}}{{\langle}v_{1}^{*}{\rangle}^{2}}$. Estimate (2.21) is proved identically, using the expression $v_{1}^{*}=v^{*}+|u|\sigma$ instead. ∎
@@
(2.22) $\int_{{\mathbb{S}}^{2}}\frac{1}{(1-\lambda|\widehat{\eta}\cdot\sigma|)^{l/2}}\,d\sigma\lesssim\frac{1}{(l-2)\lambda}(1-\lambda)^{1-l/2}.$
@@
$\int_{{\mathbb{S}}^{2}}\frac{1}{(1-\lambda|\widehat{\eta}\cdot\sigma|)^{l/2}}\,d\sigma$ $\approx\int_{0}^{1}\frac{1}{(1-\lambda x)^{l/2}}\,dx=\frac{1}{\lambda}\int_{0}^{\lambda}\frac{1}{(1-z)^{l/2}}\,dz\leq\frac{1}{(l-2)\lambda}(1-\lambda)^{1-l/2},$
@@
We are now in the position to prove the main cancellation estimate:
@@
Let $l>2$ and $1\leq p<\infty$. For $g\in C_{c}({\mathbb{R}}^{3})$, we denote
(2.23) $\mathcal{J}_{l,p}^{0}[g](v)$ $:=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\frac{{\langle}v^{*}{\rangle}^{2}|\widehat{R}_{\sigma}^{+}(u)|^{2}}{{\langle}v_{1}^{*}{\rangle}^{l}}|g(v^{*})|^{p}\mathds{1}_{|v^{*}|>2}\,d\sigma\,dv_{1}$
(2.24) $\mathcal{J}_{l,p}^{1}[g](v)$ $:=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\frac{{\langle}v_{1}^{*}{\rangle}^{2}|\widehat{R}_{\sigma}^{-}(u)|^{2}}{{\langle}v^{*}{\rangle}^{l}}|g(v_{1}^{*})|^{p}\mathds{1}_{|v_{1}^{*}|>2}\,d\sigma\,dv_{1}.$
Then there hold the estimates
(2.25) $\|\mathcal{J}_{l,p}^{0}[g]\|_{L_{v}^{\infty}}$ $\lesssim\|g\|_{L_{v}^{p}}^{p}.$
(2.26) $\|\mathcal{J}_{l,p}^{1}[g]\|_{L_{v}^{\infty}}$ $\lesssim\|g\|_{L_{v}^{p}}^{p}.$
@@
We prove estimate (2.26) as estimate (2.25) follows in an entirely analogous way. Throughout the proof, given $a\in{\mathbb{R}}^{3}$ with $a\neq 0$, we will denote $\lambda_{a}:=\frac{|a|^{2}}{{\langle}a{\rangle}^{2}}\in(0,1)$. Consider $g\in C_{c}({\mathbb{R}}^{3})$ and fix $v\in{\mathbb{R}}^{3}$. We use (2.20), (2.7) and the substitution $y:=u=v-v_{1}$ to bound
$\mathcal{J}_{l,p}^{1}[g](v)\lesssim\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\frac{{\langle}v_{1}^{*}{\rangle}^{2}|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|}{{\langle}v_{1}^{*}{\rangle}^{l}\left(1-\lambda_{v_{1}^{*}}|\widehat{v_{1}^{*}}\cdot\sigma|\right)^{l/2}}|g(v_{1}^{*})|^{p}\mathds{1}_{|v_{1}^{*}|>2}\,d\sigma\,dv_{1}$
$=\int_{{\mathbb{S}}^{2}}\int_{{\mathbb{R}}^{3}}\frac{|g(v-R_{\sigma}^{-}(u))|^{p}|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2}}{\left\langle v-R_{\sigma}^{-}(u)\right\rangle^{l-2}\left(1-\lambda_{v-R_{\sigma}^{+}(u)}|\widehat{(v-R_{\sigma}^{-}(u))}\cdot\sigma|\right)^{l/2}}\mathds{1}_{|v-R_{\sigma}^{-}(u)|>2}\,dv_{1}\,d\sigma$
$=\int_{{\mathbb{S}}^{2}}\int_{{\mathbb{R}}^{3}}\frac{|g(v-R_{\sigma}^{-}(y))|^{p}|\widehat{R}_{\sigma}^{-}(y)\cdot\sigma|}{{\langle}v-R_{\sigma}^{-}(y){\rangle}^{l-2}\left(1-\lambda_{v-R_{\sigma}^{-}(y)}|\widehat{(v-R_{\sigma}^{-}(y))}\cdot\sigma|\right)^{l/2}}\mathds{1}_{|v-R_{\sigma}^{-}(y)|^{2}>2}\,dy\,d\sigma.$
Using Proposition 2.1 to substitute $\nu:=R_{\sigma}^{-}(y)$, we obtain
$\mathcal{J}_{l,p}^{1}[g](v)\lesssim\int_{{\mathbb{S}}^{2}}\int_{{\mathbb{R}}^{3}}{\langle}v-\nu{\rangle}^{2-l}|g(v-\nu)|^{p}\mathds{1}_{|v-\nu|>2}\frac{1}{\left(1-\lambda_{v-\nu}|\widehat{(v-\nu)}\cdot\sigma|\right)^{l/2}}\,d\nu\,d\sigma$
$=\int_{{\mathbb{R}}^{3}}{\langle}v-\nu{\rangle}^{2-l}|g(v-\nu)|^{p}\mathds{1}_{|v-\nu|>2}\int_{{\mathbb{S}}^{2}}\frac{1}{\left(1-\lambda_{v-\nu}|\widehat{(v-\nu)}\cdot\sigma|\right)^{l/2}}\,d\sigma\,d\nu.$